OpenStudy (anonymous):

Radio stations are usually identified bytheir frequency. One radio station in the middle of the FM band hasa frequency of 99.0 MHz. What is itswavelength?

OpenStudy (anonymous):

When dealing with mHz, the short formula is 300/F in mHz. 300/99

OpenStudy (jamesj):

If a wave has wavelength \( \lambda \) and frequency \( f \), then its speed \( v \) is given by \[ v = f \lambda \] You can see that because by definition of frequency, in one second \( f \) entire waves go by. How long are those waves? \( \lambda \). Hence in one second the wave travels \( f \lambda \). Now, a radio wave travels at the speed of light, so \( v = c = 3.00 \times 10^{10} \ m/s\). Hence given your frequency of \( f = 99 \times 10^3 \ Hz \), the wavelength is \[ \lambda = c/f \]

OpenStudy (radar):

The given frequency is not;\[f=99\times10^{3}\]but is given as\[f =99\times10^{6}\]Hz of course/